OpenStudy (anonymous):

What is the value of the x variable in 2x + 3y = 4 x - 2y = -5

OpenStudy (anonymous):

I think you can solve it. Just find the way to eliminate y what are the number to be multiplied

OpenStudy (anonymous):

So 2(2y - 5) + 3y = 4 4y - 10 + 3y = 4 7y = 14 y = 2 x = 2*2 - 5 = -1

OpenStudy (anonymous):

you will see them they are right infront of y s

OpenStudy (anonymous):

yes you got it right

OpenStudy (anonymous):

thnx (:

OpenStudy (anonymous):

yw :)